OpenStudy (anonymous):

Can someone please explain how to solve this? sqrt[3]{sqrt[4]{27^{2}}} Preferably step by step, please! Thank you! :)

OpenStudy (anonymous):

it looks like this \[\sqrt[3]{\sqrt[4]{27^{2}}}\] Thank you! :D

OpenStudy (dumbcow):

rewrite as fraction exponent =(27^2/4)^1/3 =(27^1/2)^1/3 =27^1/6 27 = 3^3 -> (3^3)^1/6 =3^1/2 =sqrt(3)

OpenStudy (anonymous):

I'm slow sorry

OpenStudy (anonymous):

Thank you!